A class contributed $11.40 in dimes and quarters to a welfare fund. In all, there were 60 coins. How many coins of each kind were contributed?
Answer by fractalier(6550)

Call the amounts of each coin, d and q. Then we have
d + q = 60
The value equation is
10d + 25q = 1140 (we work in cents)
Now multiply the top one by ten and subtract and we get
10d + 25q = 1140
-(10d + 10q = 600)
-------------------
15q = 540
q = 36 quarters
d = 60-36 = 24 dimes